a company has found that the daily demand for its product varies inversely as the price of the product. when the price is $24, the demand is 1,000 units. what is the demand if the price is increased to $30? set up a variation problem and find the value of k. then solve for the demand when the price is $30
Answer by stanbon(75887)

a company has found that the daily demand for its product varies inversely as the price of the product.
d = k/p
---
Find "k":
when the price is $24, the demand is 1,000 units.
1000 = k/24
k = 24000
------------
So the relation is d = 24000/p
-----------------------------------
what is the demand if the price is increased to $30?
d = 24000/30
d = 800
====================
Cheers,
Stan H.
=======================
set up a variation problem and find the value of k. then solve for the demand when the price is $30
0 solutions